OpenStudy (anonymous):

Solve the equation for x. 6x-12=3x+15

OpenStudy (anonymous):

\[6x - 12 = 3x + 15\]

OpenStudy (anonymous):

subtract 3x from both sides then you get \[6x - 3x - 12 = 15\]

OpenStudy (anonymous):

add 12 to both sides so you get \[6x - 3x = 27\]

OpenStudy (anonymous):

subtrat 3x from 6x so you get \[3x = 27\]

OpenStudy (anonymous):

divide by 3 on both sides so you get \[x=9\]

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

Yes, Sorry! Thank you for explaining It!